The zinc within a copper-plated penny will dissolve in hydrochloric acid if the copper coating is filed down in several spots (so that the hydrochloric acid can get to the zinc). The reaction between the acid and the zinc is 2H+(aq)+Zn(s)→H2(g)+Zn2+(aq)2H+(aq)+Zn(s)→H2(g)+Zn2+(aq).
When the zinc in a certain penny dissolves, the total volume of gas collected over water at 25 ∘C∘C was 0.961  L L at a total pressure of 0.995  bar bar .

What mass of hydrogen gas was collected? The vapour pressure of water at 25 ∘C∘C is 0.03171 bar.
Answer #1

Solution-

The equation can be written as

2H+(aq) + Zn(s) ------> H2(g) + Zn2+(aq)

pressure = 743 - 23.78

= 719.22 mm og Hg

So the pressure = (719.22 / 760)

= 0.9463 atm

PV = nRT

=>0.9463 * 0.961 = n * 0.0821 * 298

=>0.9093/24.4658=n

=>n = 0.037 moles

So the mass of H2 = 0.037 * 2

=0.074 g